# Regge trajectories for the hidden bottom tetraquarks in the diquark picture

### v1

@@ Appendix C Determination of $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ for the $\lambda$-modes
In this section, we determine the values of $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ for the $\lambda$-modes of the hidden bottom tetraquarks. Apply Eq. (II.4) to fit the Regge trajectories for the doubly heavy mesons, the $\lambda$-modes of heavy-heavy baryons composed of one heavy quark and one doubly heavy (or one heavy-light) diquark and the $\lambda$-modes of heavy-heavy tetraquarks composed of one doubly heavy (or heavy-light) diquark and one doubly heavy (or heavy-light) antidiquark. The quality of a fit is measured by the quantity $\chi^{2}$ defined by Sonnenschein:2014jwa
@@
The masses used are listed in Tables 8 and 9. For the mesons and baryons, the listed data are the spin averaged masses. If the corresponding states are experimentally determined, the PDG data from Ref. pdg:2024 are used. For the undetermined states, the theoretical data from Ref. Ebert:2011jc ; Faustov:2021qqf are used. (In Ref. Ebert:2011jc , the mass of the $\Omega_{cbb}$ $(7/2)^{+}$(1D) state is not provided, so we use the value $11.807$ ${{\rm GeV}}$.) The theoretical data for the tetraquarks are from Faustov:2021hjs ; Faustov:2022mvs . Beside the bound states’ masses, some parameters are listed in Table 2. The masses of the axial vector diquarks are $m_{\{cc\}}=3.14$ ${{\rm GeV}}$, $m_{\{bb\}}=9.63$ ${{\rm GeV}}$ Feng:2023txx . The masses of the scalar diquarks $[cs]$ and $[bs]$ are calculated by using the parameters in Table 2 and Eq. (II.4) or (II.5) Chen:2023cws .
@@
According to Eq. (II.5), $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ are needed to determine a Regge trajectory as $m_{R_{\lambda}}$ can be calculated by using Eq. (II.5) and parameters in Table 2. Two or more states on the Regge trajectory are needed to obtain $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$. Since the hidden bottom tetraquarks are not found experimentally, we choose $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ by fitting these parameters from other systems. By fitting the data in Table 10, we have
@@
where $\mu_{\lambda}$ is the reduced masses, see Eq. (7) or (II.5). As more experimental data or more theoretical data become available, the fitted formulas will be refined.
@@
Figure 4: The radial and orbital Regge trajectories (the black lines) for mesons, baryons and tetraquarks. The dots are the spin averaged masses of mesons and baryons and the masses of tetraquarks. The used data are listed in Tables 8, 9 and 10.
Figure 5: The fitted $c_{fx_{\lambda}}$ (fig. 5) and $c_{0x_{\lambda}}$ (fig. 5), ($x_{\lambda}=L,\,N_{r}$). The data are listed in Table 10. $\mu_{\lambda}$ is the reduced masses, see Eq. (7) or (II.5). The black filled circles (mesons), the red filled squares (tetraquarks), and the green filled triangles (baryons) are the fitted $c_{fL}$ or $c_{0L}$. The black empty circles (mesons), the red empty squares (tetraquarks), and the green empty triangles (baryons) are the fitted $c_{fN_{r}}$ or $c_{0N_{r}}$. The black lines ($c_{fL}$ and $c_{0L}$) and the dashed lines ($c_{fN_{r}}$ and $c_{0N_{r}}$) are the linear fit, see Eqs. (29) and (30).

### v4

@@ Appendix D Determination of $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ for the $\lambda$-modes
In this section, we determine the values of $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ for the $\lambda$-modes of the tetraquarks with hidden bottom and charm. Apply Eq. (C.2) to fit the Regge trajectories for the doubly heavy mesons, the $\lambda$-modes of heavy-heavy baryons composed of one heavy quark and one doubly heavy (or one heavy-light) diquark and the $\lambda$-modes of heavy-heavy tetraquarks composed of one doubly heavy (or heavy-light) diquark and one doubly heavy (or heavy-light) antidiquark. The quality of a fit is measured by the quantity $\chi^{2}$ defined by
@@
The masses used are listed in Tables 13 and 14. For the mesons and baryons, the listed data are the spin averaged masses. If the corresponding states are experimentally determined, the PDG data from Ref. pdg:2024 are used. For the undetermined states, the theoretical data from Ref. Ebert:2011jc ; Faustov:2021qqf are used. (In Ref. Ebert:2011jc , the mass of the $\Omega_{cbb}$ $(7/2)^{+}$(1D) state is not provided, so we use the value $11.807$ ${{\rm GeV}}$.) The theoretical data for the tetraquarks are from Faustov:2021hjs ; Faustov:2022mvs . Beside the bound states’ masses, some parameters are listed in Table 1. The masses of the axial vector diquarks are $m_{\{cc\}}=3.14$ ${{\rm GeV}}$, $m_{\{bb\}}=9.63$ ${{\rm GeV}}$ Feng:2023txx . The masses of the scalar diquarks $[cs]$ and $[bs]$ are calculated by using the parameters in Table 1 and Eq. (C.2) or (C.3) Chen:2023cws .
@@
According to Eq. (C.3), $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ are needed to determine a Regge trajectory as $m_{R_{\lambda}}$ can be calculated by using Eq. (C.3) and parameters in Table 1. Two or more states on the Regge trajectory are needed to obtain $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$. We choose $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ by fitting these parameters from other systems. By fitting the data in Table 15, we have
@@
where $\mu_{\lambda}$ is the reduced masses, see Eq. (10) or (C.3). Eqs. (42) and (43) are obtained when $\mu_{\lambda}<3.83$ ${{\rm GeV}}$. And the formulas for the $\mu_{\lambda}>3.83$ ${{\rm GeV}}$ are not considered here. As more experimental data or more theoretical data become available, the fitted formulas will be refined.
@@
Figure 6: The radial and orbital Regge trajectories (the black lines) for mesons, baryons and tetraquarks. The dots are the spin averaged masses of mesons and baryons and the masses of tetraquarks. $N_{r}$ and $L$ are the radial and orbital quantum numbers for the $\lambda$-mode, respectively. The used data are listed in Tables 13-15.
Figure 7: The fitted $c_{fx_{\lambda}}$ [Fig. 7] and $c_{0x_{\lambda}}$ [Fig. 7], ($x_{\lambda}=L,\,N_{r}$). The data are listed in Table 15. $\mu_{\lambda}$ is the reduced masses, see Eq. (10) or (C.3). The black filled circles (mesons), the red filled squares (tetraquarks), and the green filled triangles (baryons) are the fitted $c_{fL}$ or $c_{0L}$. The black empty circles (mesons), the red empty squares (tetraquarks), and the green empty triangles (baryons) are the fitted $c_{fN_{r}}$ or $c_{0N_{r}}$. The black lines ($c_{fL}$ and $c_{0L}$) and the dashed lines ($c_{fN_{r}}$ and $c_{0N_{r}}$) are the linear fit, see Eqs. (42) and (43).